Circular No. 9038

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 F7 (LINEAR) On Mar. 31, the LINEAR team reported a single-night candidate for

the recovery of comet P/2003 H4 (cf. IAUCs 8127, 8135). Today, G. Sostero, E. Prosperi, E. Guido, and P. Camilleri reported their recovery of this comet on 60-s unfiltered CCD exposures taken remotely with a 0.35-m f/7 “Skylive” reflector at the Grove Creek Observatory, Trunkey, New South Wales; thirty co-added images show the object to have a central condensation and a diffuse coma nearly 15′′ in diameter. The available astrometry is given below. The correction to the orbital elements on MPC 56804 (and to those in the 2008/2009 Comet Handbook, p. 97) is ΔT = −0.13 day.

2009 UT α_{2000} δ_{2000} Mag. Observer
Mar. 31.30214 12^{h}53^{m}01.^{s}79 −27^{o}12′26′′.6 19.4 LINEAR
31.31444 12 53 01.53 −27 12 19.1 19.6  _{”}
31.32707 12 53 01.19 −27 12 07.6 19.8  _{”}
31.33954 12 53 00.84 −27 11 59.9 19.7  _{”}
Apr. 15.56030 12 48 07.64 −23 06 02.0 19.5 Sostero
15.57870 12 48 07.29 −23 05 40.6 19.5  _{”}
15.59400 12 48 07.00 −23 05 22.6 19.5  _{”}
15.61201 12 48 06.66 −23 05 00.8 19.7  _{”}
15.62125 12 48 06.34 −23 04 50.5 19.7  _{”}
15.63049 12 48 06.10 −23 04 39.8 19.6  _{”}

The following orbital elements by B. G. Marsden are from 231 observations, 2003 Apr. 29–2009 Apr. 15 (mean residual 0′′.6).

Epoch = 2009 June 18.0 TT

T = 2009 June 22.28618 TT  ω = 10.^{o}60506
e = 0.4902703  Ω = 226.74365 }2000.0
q = 1.7014372 AU  i = 18.15146

a = 3.3379202 AU n^{o} = 0.16161815 P = 6.10 years Additional orbital elements, residuals, and an ephemeris appear on

CBET 1767.

COMET 217P/2009 F3 (LINEAR) Comet P/2009 F3 = P/2001 MD_{7} (cf. IAUC 9031) has been assigned

the permanent number 217P (cf. MPC 65561).

2009 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green